OpenStudy (anonymous):

x sqrt(8x^3) + x sqrt(2) + 4 sqrt(2x) If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (saifoo.khan):

\[Megan-Foxx.\]

OpenStudy (saifoo.khan):

\[\Large x \sqrt{8x^3}+x \sqrt{2}+4\sqrt{2x}\]Like that?

OpenStudy (anonymous):

yeahh

OpenStudy (saifoo.khan):

\[\sqrt{2x} (2 x^2+\sqrt{x}+4) \]

OpenStudy (saifoo.khan):

something like that.

OpenStudy (anonymous):

oh okayy